A fireman is sliding down with the help of a rope. If the breaking strength of the rope is 2mg/3 where m is the mass of man, the acceleration with which the fireman should slide , so that the rope does not break ,is (1) g/2 (2) g/3(3) g/4(4) 2g/3

Grade:12th pass

1 Answers

Somi
7 years ago
2.g/3 mg-T=ma(since,sliding down)Here breaking strength is the tension of the rope i.e.=2mg/3mg-2mg/3=mamg/3=ma•°•a=g/3: